OpenStudy (anonymous):

How to find the lim as x->0 of (1-cos5x)/(4x) without hopital. What methods are available? I'm sure I could graph it, table of values it. Is there an algebra or derivative solution?

OpenStudy (anonymous):

granted that lim{x->0} (1 - cosx)/x = 0 (5/4) lim{x-0} (1 - cos(5x)) / (5x) (5/4) (0) = 0